A point starts moving in a straight line with a certain acceleration. At a time t after beginning of motion the acceleration suddenly becomes retardation of the same value.The time in which the point returns to the initial point is: 
 
   1) root2 t     2) (2 + root2) t    3) t/(root2)   4) Cannot be predicted
 

Grade:11

1 Answers

DR STRANGE
6 years ago
the ans is 1) root2 t
s= ut + at2/2
let acceleration a = p
initially the velocity 0
distance d travelled in time t
d=pt2/2............(distance travelled in time t)
after time t acceleration becomes retardation i.e. a = -p
it will travel same distance d before its veloity becomes due to retardation.
therefore distance travelled is d+d= 2d
now the particle will retrace the path backward
again 
s= ut + at2/2
2d = 0t + pt’2/2
2pt2/2 = 0t + pt’2/2
t’ = root2 t
Approve if you are satisfied :)